OpenStudy (anonymous):

Twenty-six is what percent of 40 plz help with an explanation so i can understand

OpenStudy (greatlife44):

Well, we can set up a ratio and the new can cross multiply \[\frac{ 26}{ 40 } = \frac{ x }{ 100 }\]

OpenStudy (greatlife44):

40x = 2600

OpenStudy (phi):

In English , when we say for example 1 is ½ of 2 that means (in math) 1 = ½ * 2 the "is" turns into an = sign the ½ is multiplied times the 2

OpenStudy (phi):

Twenty-six is what percent of 40 becomes 26 = x% * 40 notice the "is" turns into = the what percent (an unknown amount we call x) of 40 turns into x% * 40 also percent (%) means divide by 100 so we can change x% to x/100 finally, putting it together: \[ 26 = \frac{x}{100} \cdot 40 \] or if you like \[ \frac{40}{100} x = 26 \\ \frac{2}{5} x = 26 \\ \] to find x we multiply both sides by 5/2 to get \[ x= \frac{5}{2}\cdot 26\]

OpenStudy (phi):

the other way is to think great life did, and use ratios that is a good way if you understand it.